OpenStudy (anonymous):

Please help me.. log27 x + log27 (x+8) = 2/3

OpenStudy (john_es):

I think you mean, \[\log_{27}x+\log_{27}(x+8)=2/3\] You should first use the property, \[\log_ax+\log_ay=\log_a(x\cdot y)\] And then use the fact, \[\log_ax=z\Rightarrow x=a^z\]

OpenStudy (anonymous):

thank you but i was hoping someone could should me step by step as i do not fully understand this subject

OpenStudy (john_es):

Ok, in case you don't found anybody, I think the two properties I wrote are good points to start. When you learn them, you will be able to completely solve the problem. ;)